# MaxMin-RLHF: Towards Equitable Alignment of Large Language Models with Diverse Human Preferences

### v1

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) aligns language models to human preferences by employing a singular reward model derived from preference data. However, such an approach overlooks the rich diversity of human preferences inherent in data collected from multiple users. In this work, we first derive an impossibility result of alignment with single reward RLHF, thereby highlighting its insufficiency in representing diverse human preferences. To provide an equitable solution to the problem, we learn a mixture of preference distributions via an expectation-maximization algorithm and propose a MaxMin alignment objective for policy learning inspired by the Egalitarian principle in social choice theory to better represent diverse human preferences. We elucidate the connection of our proposed approach to distributionally robust optimization and general utility RL, thereby highlighting the generality and robustness of our proposed solution. We present comprehensive experimental results on small-scale (GPT-2) and large-scale language models (with Tulu2-7B) and show the efficacy of the proposed approach in the presence of diversity among human preferences. Our algorithm achieves an average improvement of more than 16% in win-rates over conventional RLHF algorithms and improves the win-rate (accuracy) for minority groups by over 33% without compromising the performance of majority groups, showcasing the robustness and fairness of our approach. We remark that our findings in this work are not only limited to language models but also extend to reinforcement learning in general.
@@ 1 Introduction
The alignment problem, central to developing and fine-tuning current large language models (LLMs), represents a crucial challenge in artificial intelligence, especially in ensuring these models operate in harmony with human values, preferences and social welfare (Wang et al., 2023; Christian, 2020). Reinforcement learning from human feedback (RLHF) has emerged as a pivotal approach to alignment problems, specifically aligning LLM (Wang et al., 2023; Ouyang et al., 2022b; Stiennon et al., 2022a; Ouyang et al., 2022a). RLHF operates in three steps (a) supervised fine-tuning, (2) reward learning, and (3) RL fine-tuning. Step 2 learns a reward function that is expected to represent the preference feedback of the human population. However, there has been minimal emphasis on accurately representing the diversity of human preferences and the broad spectrum of user populations. As highlighted by Aroyo & Welty (2015); Aroyo et al. (2023a, b), “the notion of ‘one truth’ in crowdsourcing responses is a myth” and we need to account for the diversity in opinions and preferences.
@@
Despite the criticality, most of the latest RLHF approaches ignore the consideration of the diversity in human preference feedback by aligning the language model with a single reward function (Wang et al., 2023; Christian, 2020; Stiennon et al., 2022a; Ouyang et al., 2022a). This assumption of a single ground truth reward is restrictive and can potentially subdue the preferences or opinions of minority groups, leading to societal biases (Figure 1). To mitigate this issue, some of the recent research proposes to learn multiple reward functions, which can then be aggregated in arbitrary manners (Bakker et al., 2022; Jang et al., 2023). On the other hand, (Ovadya, 2023) adopts a consensus-based method for aggregating human representations by emphasizing specific principles (Bai et al., 2022b; Kovač et al., 2023), which might result in the under-representation of marginalized groups (Ramé et al., 2023). Another line of research focuses on the aspect of designing multi-policy strategies by fine-tuning personalized language models towards individual rewards (Jang et al., 2023; Ramé et al., 2023; Ji et al., 2023a).
@@
Some recent literature has brought attention to the challenge of aligning single utility RLHF with diverse preferences (Bakker et al., 2022; Ramé et al., 2023; Jang et al., 2023). However, a thorough understanding of how the diversity within human sub-populations influences the overall alignment objective remains elusive. Consequently, this prompts us to pose the following question:
Is the current single reward RLHF pipeline sufficient to align with diverse human preferences?
@@
In this work, we present negative results for the above question by demonstrating the impossibility of alignment using single reward RLHF (Theorem 3.3). We introduce a notion of diversity between human subpopulations due to the differences in preference distributions and establish lower bounds on the alignment performance of single reward RLHF. However, this impossibility result naturally raises another important question:
@@
What strategies can we design to efficiently align with diverse human preferences?
In response to this question, we draw inspiration from the Egalitarian rule (Sen, 2017) to provide an equitable solution to the problem and aim to maximize the social utility objective for alignment. We propose learning a mixture of preference distributions via expectation-maximization algorithm in absence of human sub-population group information. Under the mixture distributions, we introduce the notion of MaxMin RLHF, for policy learning inspired by the Egalitarian principle in social choice theory to better represent human preferences under diversity. We elucidate the connection of our proposed approach to distributionally robust optimization and general utility RL, thereby highlighting the generality and robustness of our proposed solution. We summarize our contributions as follows.
@@
(1) An impossibility result of alignment with single reward-based RLHF. We first introduce the notation of diversity as a divergence between the human sub-population preference distributions (Definition 2) and derive lower bounds on the reward model sub-optimality (Lemma 3.2) in terms of the diversity. Finally, we establish a lower bound (Theorem 3.3) on the alignment gap due to the diversity in the human preference feedback. True to our knowledge, ours is the first work to formally characterize the sub-optimality gap with conventional RLHF methods under diverse preferences.
(2) Max-Min RLHF alignment with diverse user preferences. To provide an equitable solution to the problem, we learn a mixture of preference distributions using the Expectation-Maximization (EM) algorithm (Algorithm 2) and propose a novel MaxMin-RLHF algorithm as an egalitarian strategy to align language models with social utility objectives under diverse preferences. (Algorithm 1).
(3) A comprehensive empirical study. We present a detailed empirical analysis of our proposed approach on both small (GPT-2) and large language models (Tulu-7B). Initially, we provide empirical evidence highlighting the impossibilities of alignment with single reward RLHF, followed by demonstrating the feasibility and effectiveness of MaxMin-RLHF in achieving social utility objectives. Our algorithm achieves an average improvement of more than 16% in win-rates over conventional RLHF algorithms and improves the win-rate (accuracy) for minority groups by over 33% without compromising the performance of majority groups, showcasing the robustness and fairness of our approach.
1.1 Related Works
The current RLHF approaches (Stiennon et al., 2022b; Ziegler et al., 2020; Zhu et al., 2023) involve training a reward model based on human preference feedback and then fine-tuning the language model using proximal policy optimization (PPO) (Schulman et al., 2017). Besides PPO, DPO (Direct Preference Optimization, Rafailov et al. (2023)) directly trains the large language model using human preferences without training the reward model. A self-play-based approach such as SPIN (Chen et al., 2024) is similar to DPO but has an iterative framework. However, most of the existing alignment approaches only consider the average preference by human annotators and ignore the inherent diversity among human preferences (Casper et al., 2023; Kaufmann et al., 2023). Diverse human preferences stem significantly from various factors related to social and cultural backgrounds (Aroyo et al., 2023b, a; Denton et al., 2021a). The key factors contributing to this diversity include (i) socio-demographic backgrounds (Vogels, 2021), (ii) personal bias and context subjectivity (Denton et al., 2021b; Sandri et al., 2023)), (iii) Imperfect preferences, (Sandri et al., 2023), and (iv) Linguistic ambiguity & missing context (Sandri et al., 2023; Denton et al., 2021b; Sap et al., 2022). These factors collectively underscore the complexity of aligning LLM outputs with the diverse preferences of human users, demonstrating the importance of recognizing and addressing the multifaceted nature of user feedback. Appendix B provides a detailed context of related works.
@@ 3.1 Diversity in Human Preferences
The main shortcoming of state-of-the-art alignment approaches arises from the underlying assumption that human preferences are derived from a single latent reward model $r^{*}(\textbf{y},\textbf{x})$ (cf. (2)), which fails to account for the inherent diversity among the human sub-populations (see Figure 2). As discussed in Section 1.1, one of the key reasons for the diverse human preferences is the varied socio-demographic and socio-cultural backgrounds of human sub-populations (Aroyo et al., 2023b, a). For example, population groups with diverse demographic markers such as race, ethnicity, age groups, genders, etc., have highly varied preferences as highlighted in (Aroyo et al., 2023b, a; Denton et al., 2021a). Such diversity inevitably leads to natural sub-groups of populations among humans. Modeling this diversity in preferences for the fine-tuning of language models in RLHF is crucial, which, to the best of our knowledge, is currently missing from the literature.
@@
where $p^{*}(\mathbf{y}_{1}\succ\mathbf{y}_{2}\mid\mathbf{x})$ is the probability of preferring ${\mathbf{y}}_{1}$ over ${\mathbf{y}}_{2}$ for any given pair $(\mathbf{y}_{1},\mathbf{y}_{2})$ corresponding to prompt $\mathbf{x}$. In (4), the expectation is over a finite set of humans $h\in{\mathcal{H}}$. We next introduce the concept of human subpopulations as a hidden random variable, denoted as $u$ with distribution $\eta$, to account for the inherent diversity within the population. Specifically, $u$ represents the human subpopulation defined over a finite discrete set $\mathcal{U}:=\{\mathcal{H}_{1},\mathcal{H}_{2},\cdots,\mathcal{H}_{|\mathcal{U}|}\}$, such that $\mathcal{H}=\bigcup_{u=1}^{|\mathcal{U}|}\mathcal{H}_{u}$. The cardinality of the set $\mathcal{U}$ represents the number of sub-populations/groups present in the total human population $\mathcal{H}$. Therefore, similar to (4), we can define a human-subpopulation or group-specific preference distribution for a given pair of responses $(\mathbf{y}_{1},\mathbf{y}_{2})$ and prompt $\mathbf{x}$ as
@@
$\texttt{Diverity}\ (i,j):=\text{TV}(p_{i}^{*},p^{*}_{j}),$ (6)
where TV denotes the total variation distance between two preference distributions. By utilizing the definition of sub-population groups in $\mathcal{U}$, we can express the preference in (4) as
@@ 3.2 Reward Mismatch Due to Diversity
$\underbrace{\|\phi^{*}-\phi_{u}^{*}\|}_{\textbf{Reward mismatch}}\geq\frac{1}{2D}\cdot{\sum_{k=1}^{|\mathcal{U}|}\eta(k)\cdot\texttt{Diverity}\ (u,k)},$
@@
where $\eta$ denotes the weights distribution across human sub-population groups, $D$ denotes the upper bound on the feature representation $\|\psi({\mathbf{y}},{\mathbf{x}})\|\leq D$ for all $({\mathbf{x}},{\mathbf{y}})$, and diversity as defined in Definition 2. Proof Sketch. Here we describe the proof sketch of Lemma 3.2 with a detailed proof provided in Appendix D. We begin with the definition of sub-optimality in the learned reward for a subpopulation group $u$ as $\Delta_{u}^{r}:=\hat{\phi}_{\texttt{MLE}}-\phi_{u}^{*}$ where $\hat{\phi}_{\texttt{MLE}}$ which is the approximation to the true parameter $\phi^{*}$ which minimizes the loss function in (2). However, we know in the limit of infinite data, $\hat{\phi}_{\texttt{MLE}}$ converges to $\phi^{*}$ and hence we focus on the sub-optimality gap due to diversity as $\|\phi_{u}^{*}-\phi^{*}\|$. Using the Lipschitzness of the preference probability distribution under the Bradley-Teryy preference model (derived in Lemma C in Appendix) we lower-bound the sub-optimality gap by $\frac{1}{2D}\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}})$ and finally expanding upon the definition of $p^{*}$ as shown in (3.1), we get the final result.
@@
Remark. Lemma 3.2 indicates that the current RLHF-based reward learning paradigm (Christian, 2020; Bai et al., 2022a; Rafailov et al., 2023) will suffer sub-optimality due to diversity amongst the humans, which is highly likely in practice (Aroyo et al., 2023b). Lemma 1 implies that the degree to which the learned reward parameter diverges from optimality for a given subgroup is influenced by two key factors: the distinctiveness of that subgroup’s preferences compared to all the other subgroups, and the relative weight assigned to the subgroup in the overall preference model.
3.3 An Impossibility Result of Alignment
@@
To mathematically characterize the impossibility of aligning the language model with diverse sub-population groups, let us reconsider the RL fine-tuning optimization problem, which is given by (step 3 in RLHF in Section 2)
$\max_{{\pi}}F_{{\phi}}(\pi),$ (10)
@@
Figure 3: (Empirical Evidence of Impossibility). This figure provides empirical evidence of the impossibility of alignment (cf. Theorem 3.3) in single reward RLHF on preference dataset presented in Figure 2. Here, the task is to align a language model to generate positive sentiment responses that are concise (of shorter token length) in nature. The top row shows the sentiment score (higher is better) distribution of the language model before alignment (left) and after alignment (right). The bottom row shows the conciseness score (lower is better) distribution of the language model before alignment (left) and after alignment (right). We note that the aligned language model (top row right) can generate highly positive sentiment sentences but completely ignores the requirement of conciseness (generates longer sentences). This is happening because the humans who prefer conciseness are in minority as described in Figure 2.
@@
where we define $F_{{\phi}}(\pi):=\mathbb{E}_{\mathbf{x}\sim\mathcal{P}}\Big{[}\mathbb{E}_{\mathbf{y}\sim\pi(\cdot~{}|~{}\mathbf{x})}\bigl{[}r_{\phi}(\mathbf{y},\mathbf{x})\bigr{]}-\beta\mathbb{D}_{\textrm{KL}}\bigl{[}\pi(\cdot~{}|~{}\mathbf{x})||\pi_{\text{ref}}(\cdot~{}|~{}\mathbf{x})\bigr{]}\Big{]}$. Let us define $\pi_{\text{RLHF}}:=\mathop{\mathrm{argmax}}_{{\pi}}F_{{\phi^{*}}}(\pi)$ where $\pi_{\text{RLHF}}$ is the optimal aligned policy with single reward RLHF. On the other hand, we define a human sub-population specific optimal policy as $\pi^{*}_{\text{u}}:=\mathop{\mathrm{argmax}}_{{\pi}}F_{{\phi^{*}_{u}}}(\pi)$, where $\pi^{*}_{\text{u}}$ is the optimal aligned policy with individual subpopulation group $\mathcal{H}_{u}$. We define the alignment gap of RLHF model $\pi_{\text{RLHF}}$ to a specific user group $\mathcal{H}_{u}$ by
@@
We note that the alignment gap defined in (11) measures the discrepancy between the reward returns by the single reward RLHF model $\pi_{\text{RLHF}}$ and the optimal model $\pi_{u}^{*}$ tailored for $\mathcal{H}_{u}$ subpopulation evaluated under true reward function $r_{u}^{*}$. Next, we present our impossibility result in Theorem 3.3. {thm}[An Impossibility Result] Let $\phi^{*}$ denotes the reward parameter, which models $p^{*}$ (cf. 1), $\phi^{*}_{u}$ denotes the human sub-population group $\mathcal{H}_{u}\in\mathcal{U}$ specific reward function to model $p_{u}^{*}$, and alignment gap is as defined in (11). Then, it holds that
@@
where $\eta$ denotes the weights distribution across human sub-population groups, $D$ denotes the upper bound on the feature representation $\|\psi({\mathbf{y}},{\mathbf{x}})\|\leq D$ for all $({\mathbf{x}},{\mathbf{y}})$, $\lambda_{\psi}$ denotes the minimum eigenvalue of the feature matrix, $\beta$ is the regularization parameter of RLHF framework, and diversity as defined in Definition 2. A detailed proof of Theorem 3.3 is provided in Appendix E. We briefly describe the proof sketch of Theorem 3.3 as follows.
@@
Remark. Theorem 3.3 shows that high subpopulation diversity inevitably leads to a greater alignment gap for the model $\pi_{\text{RLHF}}$ aligned with single reward-based RLHF framework. In summary, if a subgroup exhibits distinctive preferences or constitutes a minority with a smaller representation, the resulting model from single reward RLHF setting cannot accurately reflect the sub-population’s specific preferences. We provide empirical evidence of impossibility of alignment in Figure 6.
4 Proposed Approach: Alignment with Diverse Preferences
From the statement of Theorem 3.3, we note that it is impossible to align the language model with diverse human preferences with a single reward model based RLHF. Now the natural question arises is what can we do in practice to deal with the diverse human preferences alignment problem? This is exactly what we answer in this section. We start by proposing to learn a mixture of preference distributions as follows.

### v2

# MaxMin-RLHF: Alignment with Diverse Human Preferences

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) aligns language models to human preferences by employing a singular reward model derived from preference data. However, the single reward model overlooks the rich diversity of human preferences inherent in data collected from multiple users. In this work, we first derive an impossibility result of alignment with single reward RLHF, thereby highlighting its insufficiency in representing diverse human preferences. Next, we propose to learn a mixture of reward models via an expectation-maximization algorithm and solve a MaxMin alignment objective inspired by the Egalitarian principle in social choice theory to better honor diverse human preferences. We present comprehensive experimental results on small-scale (GPT-2) and large-scale language (with Tulu2-7B)) and show the efficacy of the proposed approach in the presence of diversity among human preferences. We remark that our findings in this work are not only limited to language models but also extend to reinforcement learning in general.
Machine Learning, ICML
[…]
@@ 1 Introduction
The alignment problem, central to developing and fine-tuning current large language models (LLMs), represents a crucial challenge in artificial intelligence, especially in ensuring these models operate in harmony with human values and preferences (Wang et al., 2023; Christian, 2020). Reinforcement learning from human feedback (RLHF) has emerged as a pivotal approach to alignment problems, specifically aligning LLM (Wang et al., 2023; Ouyang et al., 2022b; Stiennon et al., 2022a; Ouyang et al., 2022a). RLHF starts with pre-training a generative (LLM) model and subsequently fine-tuning it through supervised learning on a high-quality dataset on various downstream tasks. RLHF operates in three steps (a) supervised fine-tuning, (2) reward learning, and (3) RL fine-tuning. Step 2 learns a reward function that is expected to represent the preference feedback of the human population.
However, there has been minimal emphasis on accurately representing the diversity of human preferences and the broad spectrum of user populations. As highlighted by Aroyo & Welty (2015); Aroyo et al. (2023a, b), “the notion of ‘one truth’ in crowdsourcing responses is a myth” and we need to account for the diversity in opinions and preferences.
@@
Despite the criticality, most of the latest RLHF approaches ignore the consideration of the diversity in human preference feedback by aligning the language model with a single reward (Wang et al., 2023; Christian, 2020; Stiennon et al., 2022a; Ouyang et al., 2022a). The assumption of a single ground truth reward is restrictive and can potentially subdue the preferences or opinions of minority groups, leading to societal biases (Figure 1). To mitigate this issue, some of the recent research proposes to learn multiple reward functions, which can then be aggregated in arbitrary manners (Bakker et al., 2022). On the other hand, (Ovadya, 2023) adopts a consensus-based method for aggregating human representations by emphasizing specific principles (Bai et al., 2022b; Kovač et al., 2023), which might result in the under-representation of marginalized groups (Ramé et al., 2023). Another line of research focuses on the aspect of designing multi-policy strategies by fine-tuning personalized language models towards individual rewards (Jang et al., 2023; Ramé et al., 2023; Ji et al., 2023a).
@@
As mentioned above, the recent literature has brought attention to the challenge of aligning single utility RLHF with diverse preferences. However, a thorough understanding of how the diversity within human sub-populations influences the overall alignment objective remains elusive. Consequently, this prompts us to pose the following question: Is a single reward RLHF pipeline sufficient to align with diverse human preferences?
@@
In this work, we present negative results for the above question in this work by demonstrating the impossibility of alignment using single reward RLHF (Theorem 3.3). We introduce a notion of diversity between human subpopulations due to the differences in preference distributions and establish lower bounds on the alignment performance of single reward RLHF. However, this impossibility result naturally raises another important question:
@@
What strategies can we design (or what methods can we adopt) to align with diverse human preferences?
In response to this question, we draw inspiration from the Egalitarian rule (Sen, 2017) and aim to maximize the social utility objective for alignment. We summarize our contributions as follows.
@@
(1) An impossibility result of alignment with single reward-based RLHF. We first introduce the notation of diversity (Definition 2) and then derive lower bounds on the reward model suboptimality (Lemma 3.2) in terms of diversity in human sub-population preference distributions. Finally, we establish a lower bound (Theorem 3.3) on the alignment gap due to the diversity in the human preference feedback. True to our knowledge, our work is the first to report such a result in the RLHF literature.
(2) Max-Min RLHF alignment with diverse user preferences. We propose to learn a mixture of preference distributions through the application of multiple reward functions using the Expectation-Maximization (EM) algorithm (Algorithm 2). Upon obtaining multiple reward functions specific to different human sub-populations, we introduce the MaxMin-RLHF algorithm as a strategy to align language models with social utility objectives (Algorithm 1).
(3) A comprehensive empirical study. We present a detailed empirical analysis of our proposed concepts on two language models: GPT-2 and Tulu-7B. Initially, we provide empirical evidence highlighting the impossibilities of alignment with single reward RLHF, followed by demonstrating the feasibility and effectiveness of MaxMin-RLHF in achieving social utility objectives. Our approach outperforms existing methodologies, showcasing significant performance improvements.
@@ 3.1 Diversity in Human Preferences
The main shortcoming of state-of-the-art alignment approaches arises from the underlying assumption that human preferences are derived from a single latent reward model $r^{*}(\textbf{y},\textbf{x})$ (cf. (2)), which fails to account for the inherent diversity among the human sub-populations (see Figure 2). One of the key reasons (discussed in Appendix B in ) for the diverse human preferences is the varied socio-demographic and socio-cultural backgrounds of human sub-populations (Aroyo et al., 2023b, a). For example, population groups with diverse demographic markers such as race, ethnicity, age groups, genders, etc., have highly varied preferences as highlighted in (Aroyo et al., 2023b, a; Denton et al., 2021a). Such diversity inevitably leads to natural sub-groups of populations among humans. Modeling this diversity in preferences for the fine-tuning of language models in RLHF is crucial, which, to the best of our knowledge, is currently missing from the literature.
@@
where $p^{*}(\mathbf{y}_{1}\succ\mathbf{y}_{2}\mid\mathbf{x})$ is the probability of preferring ${\mathbf{y}}_{1}$ over ${\mathbf{y}}_{2}$ for any given pair $(\mathbf{y}_{1},\mathbf{y}_{2})$ corresponding to prompt $\mathbf{x}$. In (4), the expectation is over a finite set of humans $h\in{\mathcal{H}}$. We next introduce the concept of human subpopulations as a hidden random variable, denoted as $u$ with distribution $\eta$, to account for the inherent diversity within the population. Specifically, $u$ represents the human subpopulation defined over a finite discrete set $\mathcal{U}:=\{\mathcal{H}_{1},\mathcal{H}_{2},\cdots,\mathcal{H}_{|\mathcal{U}|}\}$, such that $\mathcal{H}=\bigcup_{u=1}^{|\mathcal{U}|}\mathcal{H}_{u}$. The cardinality of the set $\mathcal{U}$ represents the number of sub-populations/groups present in the total human population $\mathcal{H}$.
Therefore, similar to (4), we can define a human-subpopulation or group-specific preference distribution for a given pair of responses $(\mathbf{y}_{1},\mathbf{y}_{2})$ and prompt $\mathbf{x}$ as
@@
$\texttt{Diversity}\ (i,j):=\text{TV}(p_{i}^{*},p^{*}_{j}),$ (6)
where TV denotes the total variation distance between two preference distributions.
By utilizing the definition of sub-population groups in $\mathcal{U}$, we can express the preference in (4) as
@@ 3.2 Reward Mismatch Due to Diversity
$\underbrace{\|\phi^{*}-\phi_{u}^{*}\|}_{\textbf{Reward mismatch}}\geq\frac{\epsilon(1-\eta(u))}{4D},$ (10)
@@
where $\epsilon:=\texttt{Diverity}(u,j)-\max_{k\neq u}\texttt{Diverity}(k,j)>0$, $D$ denotes the upper bound on the feature representation $\|\psi({\mathbf{y}},{\mathbf{x}})\|\leq D$ for all $({\mathbf{x}},{\mathbf{y}})$, and diversity as defined in Definition 2. Proof Sketch. Here we describe the proof sketch of Lemma 3.2 with a detailed proof provided in Appendix D. We begin with the definition of sub-optimality in the learned reward for a subpopulation group $u$ as $\Delta_{u}^{r}:=\hat{\phi}_{\texttt{MLE}}-\phi_{u}^{*}$ where $\hat{\phi}_{\texttt{MLE}}$ which is the approximation to the true parameter $\phi^{*}$. However, we know in the limit of infinite data under appropriate regulatory conditions, $\hat{\phi}_{\texttt{MLE}}$ converges to $\phi^{*}$, and hence we focus on the sub-optimality gap due to diversity as $\|\phi_{u}^{*}-\phi^{*}\|$. Using the Lipschitzness of the preference probability distribution under the Bradley-Tery preference model (derived in Lemma C in Appendix) we lower-bound the sub-optimality gap and finally expanding upon the definition of $p^{*}$ as shown in (3.1), we get the final result.
@@
Remark. Lemma 3.2 indicates that the current RLHF-based reward learning paradigm (Christian, 2020; Bai et al., 2022a; Rafailov et al., 2023) will suffer sub-optimality due to diversity amongst the humans, which is highly likely in practice (Aroyo et al., 2023b). Lemma 3.2 implies that the degree to which the learned reward parameter diverges from optimality for a given subgroup is influenced by two key factors: the distinctiveness of that subgroup’s preferences compared to all the other subgroups, and the relative weight assigned to the subgroup in the overall preference model.
3.3 An Impossibility Results of Alignment
@@
To mathematically characterize the impossibility of aligning the language model with diverse sub-population groups, let us reconsider the RL fine-tuning optimization problem, which is given by (step 3 in RLHF)
$\max_{{\pi}}F_{r_{\phi}}(\pi),$ (11)
@@
Figure 3: (Empirical Evidence of Impossibility). This figure validates our theoretical results in Theorem 3.3 and provides empirical evidence of the impossibility of alignment in single reward RLHF on preference dataset presented in Figure 2. Here, the task is to align the LLM to generate positive sentiment responses which are concise. We note that the aligned language model can generate highly positive sentiment sentences but completely ignores the requirement of conciseness. This is happening because the humans who prefer conciseness are in the minority as compared to humans who prefer a positive sentiment score as described in Figure 2.
@@
where we define $F_{r_{\phi}}(\pi):=\mathbb{E}_{\mathbf{x}\sim\mathcal{P}}\Big{[}\mathbb{E}_{\mathbf{y}\sim\pi(\cdot~{}|~{}\mathbf{x})}\bigl{[}r_{\phi}(\mathbf{y},\mathbf{x})\bigr{]}-\beta\mathbb{D}_{\textrm{KL}}\bigl{[}\pi(\cdot~{}|~{}\mathbf{x})||\pi_{\text{ref}}(\cdot~{}|~{}\mathbf{x})\bigr{]}\Big{]}$. Let us define $\pi^{*}_{\text{RLHF}}:=\mathop{\mathrm{argmax}}_{{\pi}}F_{r_{\phi^{*}}}(\pi)$ where $\pi^{*}_{\text{RLHF}}$ is the optimal aligned policy with single reward RLHF. On the other hand, we define a human sub-population specific optimal policy as $\pi^{*}_{\text{u}}:=\mathop{\mathrm{argmax}}_{{\pi}}F_{r_{\phi^{*}_{u}}}(\pi)$, where $\pi^{*}_{\text{u}}$ is the optimal aligned policy with individual subpopulation group $\mathcal{H}_{u}$. We define the alignment gap of RLHF model $\pi^{*}_{\text{RLHF}}$ to a specific user group $\mathcal{H}_{u}$ by
@@
We note that the alignment gap defined in (12) measures the discrepancy between the reward returns by the single reward RLHF model $\pi_{\text{RLHF}}$ and the optimal model $\pi_{u}^{*}$ tailored for $\mathcal{H}_{u}$ subpopulation evaluated under true reward function $r_{u}^{*}$. Next, we present our impossibility result in Theorem 3.3 {thm}[An Impossibility Result] Let $\phi^{*}$ denotes the reward parameter, which models $p^{*}$ (cf. 1), $\phi^{*}_{u}$ denotes the human sub-population-group $\mathcal{H}_{u}\in\mathcal{U}$ specific reward function to model $p_{u}^{*}$, and alignment gap is as defined in (12). Then, it holds that
@@
where $\epsilon:=\texttt{Diversity}(u,j)-\max_{k\neq u}\texttt{Diversity}(k,j)>0$, $\eta$ denotes the representation for the human sub-population group $u$, $D$ denotes the upper bound on the feature representation $\|\psi({\mathbf{y}},{\mathbf{x}})\|\leq D$ for all $({\mathbf{x}},{\mathbf{y}})$, $\lambda_{\psi}$ denotes the minimum eigenvalue of the feature matrix, $\beta$ is the regularization parameter of RLHF framework and diversity as defined in Definition 2. A detailed proof of Theorem 3.3 is provided in Appendix E. We briefly describe the proof sketch of Theorem 3.3 as follows.
@@
Remark. Theorem 3.3 shows that high subpopulation diversity inevitably leads to a greater alignment gap. Here, $\epsilon$ depends on the diversity among user groups, highlighting that when the diversity between a specific user group $u$ and others is significantly higher compared to inter-group diversity, it suggests that $u$ is a minority. Consequently, aligning to this particular user-group with single-reward RLHF becomes particularly challenging. Moreover, as the representation of the user-group reduces i.e $\eta(u)\to 0$, the alignment gap further increases making it harder to align to this user-group. In summary, if a subgroup exhibits distinctive preferences or constitutes a minority with a smaller representation, the resulting model from single reward RLHF setting cannot accurately reflect the sub-population’s specific preferences. We provide empirical evidence of the impossibility of alignment in Figure 5.
4 MaxMin-RLHF: One Possibility
From the statement of Theorem 3.3, it is clear that it is not possible to align diverse human preferences with a single reward RLHF. We start by noting that even if we can bypass the sub-optimality in reward learning (cf. Lemma 3.2) by learning multiple reward functions $\hat{\phi}_{u}$ for all $\mathcal{H}_{u}$, it doesn’t resolve the eventually aim of language model alignment. This is because our goal is to develop a single model $\pi^{*}$ that honors diverse user preferences without demonstrating bias towards specific groups such as minorities. To achieve that, we take motivation from the Egalitarian rule in social choice theory (Sen, 2017), which states that society should focus on maximizing the minimum utility of all individuals. Hence, we write our proposed alignment objective which maximizes the social utility as
$\pi^{*}_{\mathcal{F}}\in\arg\max_{\pi}\min_{u\in\mathcal{U}}F_{r_{\phi_{u}^{*}}}(\pi)-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi||\pi_{\text{ref}}\big{]},$ (14)
where, $F_{r_{\phi_{u}^{*}}}(\pi):=\mathbb{E}_{\mathbf{x}\sim\mathcal{P},\mathbf{y}\sim\pi(\cdot~{}|~{}\mathbf{x})}\big{[}{r_{\phi_{u}^{*}}}(\mathbf{y},\mathbf{x})\big{]}$ (cf. (3)) represents the alignment objective for the $u^{\texttt{th}}$ sub-population or group among set of humans.